40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jul 3 16:28:38 2014, Nichin, Update, Electronics, RF cables installed AS55_03-07-2014_162634.pdf
    Reply  Mon Jul 7 10:35:43 2014, Jenne, Update, Electronics, RF PDs needed 
       Reply  Mon Jul 7 17:20:09 2014, Nichin, Update, Electronics, RF PDs needed 
    Reply  Mon Jul 7 16:52:02 2014, manasa, Update, Electronics, RF cables need to be rerouted 
       Reply  Mon Jul 7 18:38:27 2014, Nichin, Update, Electronics, RF cables rerouted 1.png2.png
          Reply  Mon Jul 7 22:07:29 2014, rana, Update, Electronics, RF cables rerouted Untitled.pdf
          Reply  Tue Jul 15 12:36:17 2014, Nichin, Update, Electronics, RF cables rerouted 
Message ID: 10202     Entry time: Tue Jul 15 12:36:17 2014     In reply to: 10145
Author: Nichin 
Type: Update 
Category: Electronics 
Subject: RF cables rerouted 

Quote:

 

I have not connected them to the RF switch yet. ( until I figure out how to get both the switches working properly)

 I went into the lab and connected the RF cables to the Mux. Will take measurements for each PD henceforth.

ELOG V3.1.3-